The sum of the angles of a triangle is 180°. Find the three angles of the triangle if one angle is three times the smallest angle and the third angle is 30° greater than the smallest angle.

Answer by addingup(3677)

Let the angles be A,B, and C.
B = 3A
C = A+30
A+B+C = 180 Substitute:
A+3A+A+30 = 180
5A = 150
A = 30 and
B = 3A = 30*3 = 90
C = A+30 = 30+30 = 60